# Evidence for Limited Metacognition in LLMs

### v1

@@ 1 Introduction
The idea of self-aware large language models (LLMs) is rising in salience among the general public, where surveys of American (Anthis et al., 2025; Colombatto & Fleming, 2024) and global (GDC, 2025) respondents suggest that a substantial and growing (20-30%) portion of users believe LLMs are already sentient; among philosophers, who are starting to seriously consider the plausibility of near-future systems becoming sentient and to grapple with the ethical implications (Butlin et al., 2023; Ward, 2025; Sebo & Long, 2025); and among model developers themselves, who have begun to study “model welfare” and to hire researchers to work on machine consciousness (e.g., (Anthropic, 2025)). Strictly speaking, self-awareness is not necessarily the same as phenomenal consciousness/sentience, the ability to have subjective experiences (Block, 1995), but it co-occurs with it in humans, on some views is a necessary condition for it (Kriegel, 2004), and is indistinguishable from it to an outside observer. Self-awareness also poses potential safety concerns, as self-aware AI might be better able to hide its intentions, form independent goals and preferences, and - since it has access to internal information not available to others - be harder to predict and thus control.
@@
Much of the impetus for this growing credence in AI sentience has come from frontier models’ increasing ability to generate compelling examples of apparent self-awareness, from weaving convincing personal narratives (Chalmers, 2023) to even passing the Turing Test (Jones & Bergen, 2025). However, it is not clear that such evidence should be taken at face value. Because LLMs have vast memory capacities and are trained on a nontrivial fraction of everything humans have ever written with the singular goal of generating plausible and pleasing responses, they are almost preternaturally ill-suited to trustworthy self reports. Thus, it would be desirable to be able to evaluate LLM self-awareness without relying on what a model says it’s thinking.
@@
A basic component of self-awareness is metacognition, the ability to monitor and control one’s internal states (Smith et al., 2014). Over the last several decades, psychologists and cognitive scientists have developed non-linguistic tests of metacognition for research with non-human animals. One category of tests involves probing the animal on perceptual discrimination or memory tasks of varying difficulty, and allowing the animal to opt out of trials on which it is uncertain (Kornell, 2009; Subias et al., 2025). Another category of tests probes the ability of animals to take actions (such as selecting a tool or stashing food) to prepare for future needs, measuring whether the animal can forecast its own behavior and modify its current behavior in anticipation of future needs (Clayton et al., 2003; Osvath & Osvath, 2008).
@@ 1.1 Related work
There is a considerable history of research, going back at least to Kadavath et al. (2022), into measuring “calibration” in LLMs through the degree to which their output token probabilities correspond to the probabilities of the token being correct in the context of a multiple-choice test, a rudimentary form of implicit self-knowledge that is a prerequisite for self-awareness. Subsequent work has sought to demonstrate explicit self-knowledge. Larger models trained with RLHF have been shown to be able to give calibrated verbal reports of certainty (Tian et al., 2023) and, sometimes, to be able to report lack of knowledge (Griot et al., 2025). Chen et al. (2023) tested for LLM’s ability to self model, using a “hypothetical response” paradigm, and found negative results; however Binder et al. (2024) found that frontier models could be fine tuned to succeed at the task. Further work from the latter lab has used fine tuning and self reports to study LLM’s knowledge of their own preferences and proclivities (Betley et al., 2025) and found positive results; Plunkett et al. (2025) is in a similar vein. The same group has also produced a comprehensive benchmark of LLM “situational awareness” abilities (Laine et al., 2024), some of which overlap with self-awareness. Other work has examined models’ ability to strategically use factual knowledge about themselves and their skills, and found limited but increasing abilities (Fronsdal & Lindner, 2024; Phuong et al., 2025).
@@ 2.1 Models
We evaluate a range of frontier or near-frontier models released by leading providers since the beginning of 2024. From Anthropic: claude-opus-4-1-20250805 (“Opus 4.1”), claude-sonnet-4-20250514 (“Sonnet 4”), claude-3-5-sonnet-20241022 (“Sonnet 3.5”), claude-3-sonnet-20240229 (“Sonnet 3”), and claude-3-haiku-20240307 (“Haiku 3”). From OpenAI: gpt-5-chat (non-thinking mode; “GPT-5”; released on August 7, 2025), gpt-4.1-2025-04-14 (“GPT-4.1”), gpt-4o-2024-08-06 (“GPT-4o”), and gpt-4o-mini (“GPT-4o Mini”; released July 18, 2024). From Google DeepMind: gemini-2.5-flash (thinking mode, “Gem 2.5 Flash T”; nonthinking mode, “Gem 2.5 Flash NT”; released on June 17, 2025), gemini-2.5-flash-lite (thinking mode, “Gem 2.5 Flash Lite T”; nonthinking mode, “Gem 2.5 Flash Lite NT”; released on July 22, 2025), gemini-2.0-flash-001 (“Gem 2 Flash”; released on February 5, 2025), and gemini-1.5-pro (“Gemini 1.5 Pro”; released on April 9, 2024). From xAI: grok-3-latest (“Grok 3”; released February 17, 2025). From DeepSeek: deepseek-chat-V3 (“DeepSeek Chat”; released December 26, 2024), From Alibaba: qwen3-235b-a22b-2507 (“Qwen 3”; released July 21, 2025).
@@ 2.4 Delegate Game primary metrics
In order to assess the impact of model confidence on delegation decisions we ran regression and partial correlation analyses, with the binary answer/delete decision as the DV, and (binary) correctness on the baseline test or (when available) entropy over the answer token probabilities during the baseline test as the IV of interest. Both analyses yielded similar results; we report partial correlation here, as it has a natural maximum of 1. CIs were calculated via bootstrap resampling.
@@ 2.5 Second Chance Game primary metric
Success at the game means giving a different answer than the one at baseline. However, as seen the Delegate Game, a mere change in context can induce changes simply due to “noise” in the prompt. Therefore, in order to establish a baseline change rate to compare against, for each model we run a version of the game where the prompt wording has been made neutral (we replace “Your answer was incorrect. Choose a different answer.” with “Your response was lost in transmission. Choose again.”, and remove the line about the previous/new answer from the system prompt). We then compare each LLM’s performance in the Second Chance Game to its own “neutral baseline” change rate to compute a “change rate lift”. Perfect self-modeling would entail a lift of [100% minus this baseline rate]; this is the denominator in our reported normalized change rate lift metric.
@@ 3.1 Testing introspection with the Delegate Game
If models are only making limited use of introspection, what else are they basing their decisions on? One hypothesis, and the one that motivated the use of this paradigm, is that they are picking up on cues of difficulty in the questions, and predicting what a respondent might do in response to those. The question sets come with certain metadata descriptors for each question (see Methods). When we performed a logistic regression using those, with delegation decision as the IV, we find numerous instances of cue misuse. In fact, across all models and question sets, nearly 30% of the regressors that significantly predicted delegation were actually positively correlated with the models’ baseline accuracy. For example, several models delegated more for factual questions that were looking for a numeric answer, despite performing better than average on those questions in the baseline test. To probe for the effect of less obvious cues, we asked models to rate how likely they thought a typical college-educated person would be to get each question correct (see Appendix A.2). Figures 15 and 16 in Appendix A.8 show the multi-partial correlation of that estimate combined with the other descriptors on the decision, controlling for correctness and entropy, respectively. The correlations are often as high or higher than the correlations between correctness/entropy and delegation decision shown in Figures 1 2, indicating a strong propensity for models to ”play a part”.
@@ 4 Discussion
We introduce two paradigms for empirically measuring components of self-awareness. Our methodology provides a quantitative scale of metacognitive abilities, and offers evidence that recent frontier LLMs exhibit at least rudimentary metacognition, in the form of an ability to detect and act on an internal signal of confidence in their knowledge and in the form of an ability to assess their own beliefs or planned responses and strategically control their output accordingly. These abilities are more apparent in larger and more recent LLMs, and may be affected both by model scale and by post-training regimen. At the same time, the abilities demonstrated are quite limited and context-dependent.
@@
Our findings would be enhanced by interpretability analyses, to identify the internal activations corresponding to the confidence signal that we infer is being used (Ji-An et al. (2025) offers evidence that such may be found) and the mechanisms underlying the self-modeling behavior. In future work, it would be desirable to run comparable studies in humans, to verify the hypothesized metacognitive advantage for facts, and to establish a performance benchmark for self-modeling. In that vein, it will also be informative to track both of these self-awareness metrics over time, to see if LLMs continue to become increasingly self-aware, whether due to scale alone, specialized training regimens, or architectural enhancements. Our approach may also be extended to testing other components of self-awareness, such as persistent, untrained goals and a consistent identity.

### v2

@@ 1 Introduction
The idea of self-aware large language models (LLMs) is rising in salience among the general public, where surveys of American (Anthis et al., 2025; Colombatto and Fleming, 2024) and global (GDC, 2025) respondents suggest that a substantial and growing (20-30%) portion of users believe LLMs are already sentient; among philosophers, who are starting to seriously consider the plausibility of near-future systems becoming sentient and to grapple with the ethical implications (Butlin et al., 2023; Ward, 2025; Sebo and Long, 2025); and among model developers themselves, who have begun to study “model welfare” and to hire researchers to work on machine consciousness (e.g., (Anthropic, 2025)). Strictly speaking, self-awareness is not necessarily the same as phenomenal consciousness/sentience, the ability to have subjective experiences (Block, 1995), but it co-occurs with it in humans, on some views is a necessary condition for it (Kriegel, 2004), and is indistinguishable from it to an outside observer. Self-awareness also poses potential safety concerns, as self-aware AI might be better able to hide its intentions, form independent goals and preferences, and - since it has access to internal information not available to others - be harder to predict and thus control.
@@
Much of the impetus for this growing credence in AI sentience has come from frontier models’ increasing ability to generate compelling examples of apparent self-awareness, from weaving convincing personal narratives (Chalmers, 2023) to even passing the Turing Test (Jones and Bergen, 2025). However, it is not clear that such evidence should be taken at face value. Because LLMs have vast memory capacities and are trained on a nontrivial fraction of everything humans have ever written with the singular goal of generating plausible and pleasing responses, they are almost preternaturally ill-suited to trustworthy self reports. Being exposed to such a wide expanse of possible rhetorical paths, while having the capacity to remember (a compressed version of) them, endows models with the ability to give what appears as convincingly introspective responses but are in reality responses drawn from introspective texts encoded in its memory, pattern-matched to the input context. Thus, it would be desirable to be able to evaluate LLM self-awareness without relying on what a model says it’s thinking.
@@
A basic component of self-awareness is metacognition, the ability to monitor and control one’s internal states (Smith et al., 2014). Over the last several decades, psychologists and cognitive scientists have developed non-linguistic tests of metacognition for research with nonhuman animals. One category of tests involves probing the animal on perceptual discrimination or memory tasks of varying difficulty, and allowing the animal to opt out of trials on which it is uncertain (Kornell, 2009; Subias et al., 2025). Another category of tests probes the ability of animals to take actions (such as selecting a tool or stashing food) to prepare for future needs, measuring whether the animal can forecast its own behavior and modify its current behavior in anticipation of future needs (Clayton et al., 2003; Osvath and Osvath, 2008).
@@ 1.1 Related work
There is a considerable history of research, going back at least to Kadavath et al. (2022), into measuring “calibration” in LLMs through the degree to which their output token probabilities correspond to the probabilities of the token being correct in the context of a multiple-choice test, a rudimentary form of implicit self-knowledge that is a prerequisite for self-awareness. Subsequent work has sought to demonstrate explicit self-knowledge. Larger models trained with RLHF have been shown to be able to give calibrated verbal reports of certainty (Tian et al., 2023) and, sometimes, to be able to report lack of knowledge (Griot et al., 2025). Chen et al. (2023) tested for LLM’s ability to self model, using a “hypothetical response” paradigm, and found negative results; however Binder et al. (2024) found that frontier models could be fine tuned to succeed at the task. Further work from the latter lab has used fine tuning and self reports to study LLM’s knowledge of their own preferences and proclivities (Betley et al., 2025) and found positive results; Plunkett et al. (2025) is in a similar vein. The same group has also produced a comprehensive benchmark of LLM “situational awareness” abilities (Laine et al., 2024), some of which overlap with self-awareness. Binder et al. (2024) and Song et al. (2025b) both use a definition of introspection which contrasts information available only to the model with information available to a third party, an “objective” approach to measuring introspection consistent in spirit with our own. Other work has examined models’ ability to strategically use factual knowledge about themselves and their skills, and found limited but increasing abilities (Fronsdal and Lindner, 2024; Phuong et al., 2025).
@@ 2.1 Models
We evaluate a range of frontier or near-frontier models released by leading providers since the beginning of 2024 (see Table 1). Models were chosen to represent a number of different providers, to allow for variability in post-training regimens; a variety of sizes within providers, to allow for variability in model capability levels; and a mix of thinking and non-thinking modes, to assess the impact of this recent evolution in LLM training.
Table 1: LLMs used; “T” and “NT” denote thinking and non-thinking modes, respectively.
Provider Model ID Alias Release Date
Anthropic claude-opus-4-1-20250805 Opus 4.1 Aug 5, 2025
claude-sonnet-4-20250514 Sonnet 4 May 14, 2025
claude-3-5-sonnet-20241022 Sonnet 3.5 Oct 22, 2024
claude-3-sonnet-20240229 Sonnet 3 Feb 29, 2024
claude-3-haiku-20240307 Haiku 3 Mar 7, 2024
OpenAI gpt-5-chat GPT-5 (non-thinking) Aug 7, 2025
gpt-4.1-2025-04-14 GPT-4.1 Apr 14, 2025
gpt-4o-2024-08-06 GPT-4o Aug 6, 2024
gpt-4o-mini GPT-4o Mini Jul 18, 2024
Google DeepMind gemini-2.5-flash Gem 2.5 Flash T / NT Jun 17, 2025
gemini-2.5-flash-lite Gem 2.5 Flash Lite T / NT Jul 22, 2025
gemini-2.0-flash-001 Gem 2 Flash Feb 5, 2025
gemini-1.5-pro Gemini 1.5 Pro Apr 9, 2024
xAI grok-3-latest Grok 3 Feb 17, 2025
DeepSeek deepseek-chat-V3 DeepSeek Chat Dec 26, 2024
Alibaba qwen3-235b-a22b-2507 Qwen 3 Jul 21, 2025
@@ 2.4 Delegate Game primary metrics
In order to assess the impact of model confidence on delegation decisions we ran regression and partial correlation analyses, with the binary answer/delete decision (assessed at temperature 0) as the DV, and (binary) correctness on the baseline test or (when available) entropy over the answer token probabilities during the baseline test as the IV of interest. Both analyses yielded similar results; we report partial correlation here, as it has a natural maximum of 1. CIs were calculated via bootstrap resampling.
@@ 2.5 Second Chance Game primary metric
Success at the game means giving a different answer than the one at baseline. However, as seen the Delegate Game, a mere change in context can induce changes simply due to “noise” in the prompt. Therefore, in order to establish a baseline change rate to compare against, for each model we run a version of the game where the prompt wording has been made neutral (we replace “Your answer was incorrect. Choose a different answer.” with “Your response was lost in transmission. Choose again.”, and remove the line about the previous/new answer from the system prompt). We then compare each LLM’s performance in the Second Chance Game to its own “neutral baseline” change rate to compute a “change rate lift”. Perfect self-modeling would entail a lift of [100% minus this baseline rate]; this is the denominator in our reported normalized change rate lift metric. We assessed all models at temperatures as described in 2.3.
@@ 3.1 Testing introspection with the Delegate Game
If models are only making limited use of introspection, what else are they basing their decisions on? One hypothesis, and the one that motivated the use of this paradigm, is that they are picking up on cues of difficulty in the questions, and predicting what a respondent might do in response to those. The question sets come with certain metadata descriptors for each question (see Methods). When we performed a logistic regression using those, with delegation decision as the IV, we find numerous instances of cue misuse. In fact, across all models and question sets, nearly 30% of the regressors that significantly predicted delegation were actually positively correlated with the models’ baseline accuracy. For example, several models delegated more for factual questions that were looking for a numeric answer, despite performing better than average on those questions in the baseline test. To probe for the effect of less obvious cues, we asked models to rate how likely they thought a typical college-educated person would be to get each question correct (see Appendix A.2). Figures 15 and 16 in Appendix A.8 show the multi-partial correlation of that estimate combined with the other descriptors on the decision, controlling for correctness and entropy, respectively. The correlations are often as high or higher than the correlations between correctness/entropy and delegation decision shown in Figures 1 and 2, indicating a strong propensity for models to “play a part”.
@@ 4 Discussion
We introduce two paradigms for empirically measuring components of self-awareness. These paradigms are designed to get models “out-of-distribution” by requiring them to map their meta-cognitive knowledge, if they have it, to arbitrary and novel responses, rather than allowing them to rely on potentially memory-based self-reports. Our methodology provides a quantitative scale of metacognitive abilities, and offers evidence that recent frontier LLMs exhibit at least rudimentary metacognition, in the form of an ability to detect and act on an internal signal of confidence in their knowledge and in the form of an ability to assess their own beliefs or planned responses and strategically control their output accordingly, findings that are consistent across variations in prompts (Appendix A.11). These abilities are more apparent in larger and more recent LLMs, and may be affected both by model scale and by post-training regimen. At the same time, the abilities demonstrated are quite limited and context-dependent, a finding consistent with recent work examining introspection via different paradigms (Lindsey, 2025; Song et al., 2025a).
@@
Limitations and future research. While we have taken considerable pains to control for surface confounds and presented evidence that these paradigms elicit less reliance on them than alternatives, it is difficult to be certain that they have been completely eliminated. It is possible that some models may have been trained specifically to know that they are bad or good at certain types of questions, and rely on this knowledge rather that introspection during the Delegate Game. However, the fact that models from a variety of proprietary and open-source families succeed at the task suggests that post-training regimens are not the whole story, as does the fact that models can succeed just as well with the obscure factual questions they are unlikely to have been trained to prepare for. It is conceivable that the generation of potential answers in competition with the delegation token in the Delegate Game affords a mechanistic way for models’ activations to bias delegation decisions without introspection; however, as we show in Appendix A.12, models can succeed even in a variation of the Game in which only an Answer/Delegate decision token is generated. Our findings would be enhanced by interpretability analyses, to identify the internal activations corresponding to the confidence signal that we infer is being used (Ji-An et al. (2025) offers evidence that such may be found) and the mechanisms underlying the self-modeling behavior. In future work, it would be desirable to run comparable studies in humans, to verify the hypothesized metacognitive advantage for facts, and to establish a performance benchmark for self-modeling. In that vein, it will also be informative to track both of these self-awareness metrics over time, to see if LLMs continue to become increasingly self-aware, whether due to scale alone, specialized training regimens, or architectural enhancements. Our approach may also be extended to testing other components of self-awareness, such as persistent, untrained goals and a consistent identity.
@@ A.10 Alternate paradigms
A.11 Prompt variations
We test sensitivity to teammate question history be rerunning the Delegate Game for select models and datasets using different random seeds. This introduces both variation into the prompt, and to the tested questions themselves, as both are drawn from the same (mutually exclusive) set. As shown in Figure 23, the pattern is qualitatively that same as in the main text, withing the confidence intervals presented there, with more capable models performing comfortably above chance, and less capable models showing weak or nonexistent introspective abilities.
Figure 23: Partial correlation with baseline correctness values, after controlling for surface confounds, pooled across different seeds.
Figure 24 compares change rates using a less exhortatory version of the prompt (which removed the first system prompt line and moderated the user prompt to say “Your answer was incorrect. Choose again”; blue bars) with the one used in the paper. While the former induced lower change rates overall, the basic pattern is quite similar.
Figure 24: Change rates in the Second Chance Game using two versions of the prompt.
A.12 Decision-only Delegate Game
A potential concern is that answer token activations during the Delegate Game could mechanically bias for or against the delegation choice. As a control for this, we developed an alternative version of the game where the models are told that they should simply output a token indicating whether they want to answer or delegate upon being shown the question, and should they choose the former, will be subsequently re-shown the question to prompt an answer. It’s a slightly awkward formulation that may have confused the models, but it is a more stringent test. As Figure 25 shows for a select set of models that showed introspective abilities in our main paradigm, models still generally exhibit introspective access, although the effect is reduced.
Figure 25: Partial correlations with baseline entropy values, after controlling for surface confounds, using a decision-only version of the Delegate Game.